OpenStudy (anonymous):

I'LL GIVE YOU MEDAL IF YOU'RE A BATMAN FAN 1. Write the following expression using a single exponent. 85 ÷ 8-2 87 8-7 83 8-10

Nnesha (nnesha):

\[8^5 \div 8^{-2}\]when we divide same base , `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (theloshua):

Im a batman fan... Thats my fave superhero ^-^

OpenStudy (anonymous):

ME TOO

OpenStudy (theloshua):

i shud get a medal just for that :3

OpenStudy (anonymous):

I BEAT ALL OF THE ARKHAM GAMES IN ONE DAY

Nnesha (nnesha):

here is an example \[\huge\rm 2^{2} \div 2^{-3} = 2^{2-(-3)}=2^{2+3}=2^{5}\]

OpenStudy (theloshua):

Ive watched all the old and new movies and read countless comic books ;)

OpenStudy (anonymous):

ME TOO

OpenStudy (anonymous):

BYE THE WAY I'M A GIRL 13 YR. OLLD

OpenStudy (anonymous):

OHH I GET ITS 7

OpenStudy (anonymous):

HERE'S ANOTHER Which of the following is an equivalent representation of 3-3 ?

OpenStudy (anonymous):

A.1/9 B.1/12 C.1/12 D.1/64

Nnesha (nnesha):

well you need to know exponent rules \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] flip the fraction to change the sign of the exponent

OpenStudy (anonymous):

ummmmmmmmm. is it b.?

Nnesha (nnesha):

`write these down in ur notebook` when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\] and when we divide same base , `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

Nnesha (nnesha):

http://www.jug.net/wt/exponentrules.gif

OpenStudy (anonymous):

a.?

Nnesha (nnesha):

why A ??? :=)

OpenStudy (anonymous):

wait wait!!!!! C.?

Nnesha (nnesha):

haha why C ?

Nnesha (nnesha):

guess what guessing isn't a good option right ?

Nnesha (nnesha):

\[\huge\rm 3^{-3}\] is this ur question right ?

OpenStudy (anonymous):

YEAH

Nnesha (nnesha):

okay so how would you convert it to positive exponent \[3^{-3} = ????\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] flip the fraction to change the sign of the exponent \(\color{blue}{\text{End of Quote}}\) remember the exponent rule ?

OpenStudy (anonymous):

YOU WANNA KNOW SOMETHING FUNNY? IT WAS C.

OpenStudy (anonymous):

I'M LAUGHING OUT LOUD

Nnesha (nnesha):

well ur answer choices isn't correct.

OpenStudy (future_einstein):

Batman. and my little pony!!!!!

OpenStudy (future_einstein):

@TwilightSparkle2015

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @TwilightSparkle2015 A.1/9 B.1/12 C.1/12 D.1/64 \(\color{blue}{\text{End of Quote}}\) see B and C are the same :/

OpenStudy (anonymous):

MY BAD C. WAS 1/27

Nnesha (nnesha):

good. so 3^{-3} = ?

OpenStudy (anonymous):

SORRY I GOT TO GO TO CLASS NOW. THANK YOU GUYS FOR HELPING ME TODAY

Nnesha (nnesha):

okay. have a nice day.